# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ III Calculating electron-phonon coupling
We will refer to $\Pi^{\prime}$ as an auxilary self-energy since it has dimension (energy)2. Note that the overall phase associated with gauge non-invariance of $\mathfrak{g}_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ does not affect its off-diagonal elements.
@@
To understand the origin of phonon softening, we examine ab initio and many-body results based on Eq.’s(2) and (6). This requires a modification of the QE+EPW results, since the DFPT already incorporates the main phonon renormalizations (Appendix B). In order to calculate $\Pi^{\prime}$ in Eq.(8), we must invert the dynamical matrix to get a handle on the nature of the bare phonons. This involves two issues. (i) Ambiguity in inverting the dynamical matrix, which we resolve in Section VII and Appendix C. And (ii) We need the off-diagonal self-energy elements, which are essential for the adiabatic term, even though these are unimportant for the non-adiabatic corrections, extraction of off-diagonal self-energies elements is discussed in Appendix A.
IV What determines $\boldsymbol{Q}_{CDW}$?
@@
The debate here has been on whether $\boldsymbol{Q}_{CDW}$ is dominated by the susceptibility in Eq.(7) or the electron-phonon matrix elements in Eq.(3). However there is another potential mechanism. In Eq. 4 , the pre-factor contains $1/{\omega_{\boldsymbol{q}}}$ in $|\mathfrak{g}|^{2}$. This, following Hedin[Hedin], can be replaced by $1/{\Omega_{\boldsymbol{q}}}$, which acts as a Stoner denominator and causes the diagonal element $\Pi_{\lambda\lambda}$ to diverge when $\Omega_{\bf q}\to 0$ for $\boldsymbol{q}=\boldsymbol{Q}_{CDW}$ as $T\to T_{CDW}$. A similar argument has been invoked for the CDW physics in the cuprates.[Gutzcharge]
@@
Fig. 2 presents a decomposition of the self-energy of all phonon modes in panel (a) along the $\Gamma\rightarrow M$ symmetry line. We first focus on the soft-mode mode (dark green lines). $\Pi$ (panel h) shows a sharp, nearly-divergent peak at the soft-mode $\boldsymbol{q}$-vector, which arises entirely from the $1/\Omega_{q}$-factor (panel g). In contrast, $\Pi^{\prime}=\omega\Pi$ (panel d) has a weak peak at the same $\boldsymbol{q}$-vector. This sharpening of a weak peak is a typical effect of a Stoner denominator. If the electron-phonon matrix element were $\boldsymbol{q}$-independent, then $\Pi^{\prime}$ would peak at the same $\boldsymbol{q}$-value as the susceptibility $\chi$ (dashed red curve in panel b), which is not the case. Similarly, if $\chi$ were constant, the peak would follow $G$ (green curve in panel b), but that is also not the case. The product $G\chi$ (panel c) comes closer, indicating that both $\chi$ and $G$ possess significant $\boldsymbol{q}$ and $\boldsymbol{k}$ dependence, such that the full integral over $\boldsymbol{k}$ correctly reproduces the self-energy peak in panel d. The peaks in both $\Pi$ and $\Pi^{\prime}$ (solid-black vertical lines) are in reasonable agreement with the experimental value $\boldsymbol{Q}_{CDW}\simeq 2~{}{\overrightarrow{\Gamma M}}/3$ (black vertical dashed line). Note however that the off-diagonal terms in $\Pi^{\prime}$ [panel f] due to the phonon mode-mode coupling can lead to a small additional shift of the $\boldsymbol{Q}_{CDW}$.
@@
When we extend preceding discussion to encompass the the self-energy in the full Brillouin zone, we see that (Fig. 1 (b)) the largest self-energy peak falls along the $M\rightarrow K$-line. This sharp peak is associated with a weak nesting on the Fermi surface, see Appendix A for details. Insight can be gained by recalling that in many superconductors, the superconducting gap can be captured in the optical spectra only in the dirty-limit where momentum conservation breaks down[MatBar]. A similar mechanism may be at play with the second peak along the $M\rightarrow K$-line in our case and its absence in experiments. Note that Figure 1(b) is for the clean limit. To capture the dirty limit, however one must incorporate the breaking of the momentum conservation due to random impurities, which requires averaging $G_{\lambda,\lambda^{\prime}}$ over $\boldsymbol{k}$, leading to the form of Eq. 6. Fig. A.2 (b) (Appendix A) shows that the averaging procedure smears out this weak nesting feature while preserving the peak along $\Gamma M$ with a modest loss of intensity. A similar conclusion was reached in Ref. Plummer.
V Coupling of bare phonon bands
@@
To extract the bare phonons, the divergent $1/\Omega_{\lambda,\boldsymbol{q}}$-factor in Fig. 2(g) must be replaced by $1/\omega_{\lambda,\boldsymbol{q}}$ (Appendix C). Once this is done, we can go beyond the above single mode picture by analyzing the $T$-dependence of the diagonal self-energies in Figs. 1 (b) and (c). We discuss the resulting anticrossing phenomena by considering two distinct cases below.
V.1 Longitudinal-transverse anticrossing
@@
Figure 1 (a) shows that reducing the temperature (right to left) results in crossing of the modes 2 and 3 and also of modes 6 and 7 along the $\Gamma-M$ line (see Fig. 2 for notation of phonon modes). Note that the optical modes 6 and 7 display at low temperatures an unusual Mobius topology along the $\Gamma\to M\to K\to\Gamma$ line: if one follows the dispersion of mode 6, starting from the left-hand side $\Gamma$ point, one ends up on mode 7 on the right-hand side. Repeating this process into the next $\Gamma$-point by using symmetry to jump to the left-hand side $\Gamma$-point, and starting on mode 7, we see that only after winding twice about this path do we return to the original point. This behavior can be understood by noting that the longitudinal and transverse modes are orthogonal, so that we normally expect them to cross. However, in an anisotropic material, eigenmodes are strictly longitudinal/transverse only along the high symmetry lines but this is not be the case along a general direction in Brillouin zone, so that the bands then must anti-cross and exchange wavefunction character. Since modes 6 and 7 can only cross along $\Gamma-M$, they must wind twice to go from $6\rightarrow 7\rightarrow 6$ before returning to the starting mode.
@@
A similar phenomenon is at play for modes 2 and 3 in Fig.1, although it is hard to see the crossing between these modes because they are acoustic modes with vanishing frequencies at $\Gamma$. However, the most dispersive mode near $\Gamma$ is longitudinal (L, numbered 3) at the left-hand side $\Gamma$-point. The resulting mode is numbered 2 after the $\Gamma\to M\to K\to\Gamma$ loop is closed. Similar winding numbers have been observed in the cuprates under pressure[Joh], see also Ref. Zak. At higher temperatures (right column in Fig.1) both anti-crossings disappear, suggesting that they are related to phonon softening and electron-phonon interactions. Notably, crossing of bands 6 and 7 is symmetry protected only along the $\Gamma M$ line, where the crossing point forms the tip of a phononic Dirac dispersion. Since this band crossing is driven by phonon softening, it constitutes an electron-phonon-coupling induced topological phase transition.
@@
Insight into the mode crossing of modes 6 and 7 is also obtained by looking at the self-energies. As $T$ increases, the longitudinal mode 7 hardens and stops overlapping the transverse mode. At the crossing point, a part of the self-energy, especially along $K$ to $\Gamma$ jumps from the lower to the higher (now longitudinal) mode. [The strong spectral weight is always associated with the upper modes near $K$, but due to the change of connectivity, the labels of the modes change.] Effects of anticrossing can, however, still be seen in Figs. 1(b,c), where the spectral weight of the upper mode decreases substantially below the anticrossing temperature, i.e., as $T$ is lowered, the modes start to anticross, and more longitudinal character transfers to the lower modes, along with the strong electron-phonon coupling.
@@
V.2 Longitudinal-longitudinal anticrossing
Here, the anticrossing of the three longitudinal modes is harder to see in the dispersions as there are no symmetry-allowed crossing. Kohn-anomaly related anticrossings may thus be taken as a demanding test of a phonon code. Although we have shown that QE captures mode-crossing effects for modes 6 and 7, let us now consider modes 3 and 8, which are the lowest and highest longitudinal modes. As expected, the self-energy of mode 3 decreases with increasing $T$. However, the self-energy of mode 8 increases with increasing $T$, and as Figs. 1(b,c) indicate, this is due to transfer of spectral weight from mode 3 (note the anomalous peak along the $MK$ line marked by arrow in panel (b)), which shifts to mode 8 (arrow in panel (c)) with increasing temperature. These results unambiguously demonstrate that the mode softening starts in the highest longitudinal mode, and as temperature decreases the spectral weight transfers to successively lower longitudinal modes via mode coupling, with further spectral weight transfer between nominally longitudinal and transverse modes, since they are not strictly orthogonal at arbitrary $\boldsymbol{k}$-points. In the remainder of this paper, we explore the dynamical matrix to try to extract a coherent picture in terms of ‘bare’ phonon frequencies.
@@ V.2 Longitudinal-longitudinal anticrossing
Figure 3: (color online) Phonon self-energies. From left to right 0.0057, 0.009, 0.02 eV smearing. From top to bottom: (a) clean-limit and (b) dirty limit; projected in color on phonon bands (color codes from dark to light map into weak to strong phonon self-energy).
VI Dynamical matrix and bare phonon frequencies
We can obtain the bare frequencies $\omega_{\bf q}(T)$ at various temperatures by inverting the dynamical matrix. We would expect the bare frequencies to be approximately independent of $T$, but that is not the case, see Appendix B for details. Note that phonon anharmonicity[anharm] is implicitly included in the first-principles DFPT calculations.)
We must first settle on the form of the dynamical matrix. In Appendix C, we show that, despite some ambiguity, the appropriate choice is:
@@ VI Dynamical matrix and bare phonon frequencies
which provides a conventional matrix equation for the bare phonon frequencies in terms of the dressed ones. Fig. 4 shows the bare frequencies (solid lines) obtained by neglecting the off-diagonal terms in Eq. 10. Although phonon softening can be seen, the mode anticrossings, which are controlled by the off-diagonal terms are absent.
@@
Figure 4: (color online) Single mode calculations for the three longitudinal modes (modes 3, 6, and 8). Solution of Eq. 20 for $\sigma$ =0.0565 Ry, comparing the dressed ($\Omega_{q}$, dashed lines) and bare frequencies ($\omega_{q}$, solid lines).
@@
To understand why $\Omega_{\bf q}$ becomes soft, we rewrite the single mode solution of Eq. 1 as $\Omega_{\bf q}^{2}=\omega_{\bf q}^{2}(1-S_{\bf q})$, where $S_{\bf q}=-2\Pi^{\prime}(\bf q,\Omega_{\bf q})/\omega_{\bf q}^{2}$ is an effective Stoner factor[diSalvo1977]. The instability (soft phonon) occurs at that $\bf q$ for which $S_{\bf q}\rightarrow 1$ first. At $\bf Q_{CDW}$, the bare frequency is $\omega_{\bf Q_{CDW}}=\sqrt{-2\Pi^{\prime}(\bf Q_{CDW})}$. If $\omega_{\bf q}$ depends weakly on $\bf q$, then the soft mode occurs when $|\Pi^{\prime}(\bf q)|$ is at its maximum.
VI.1 Multimode calculations
@@
Figure 5: (color online) (a) Dressed (blue dashed lines) and bare (red solid lines) phonons for $\sigma$ = 0.02 Ry smearing. (b) Same as (a) but for $\sigma$ = 0.0057 Ry. (c) Temperature dependence of the bare phonons for $\sigma$ = 0.02 Ry (orange solid lines) and 0.0057 Ry (green dashed lines). (d) Temperature dependence of the dressed phonons, with colors representing a various phonon modes: 2 (orange), 3 (red), 6 (magenta), 7 (green), and 8 (blue). $\sigma$ = 0.0057 Ry (dashed lines) and 0.02 eV (dotted lines), while the solid lines ($\sigma$ = 0.02 Ry bare phonons) are our best estimate of the high-$T$ limit.
@@
Figs. 5(a) and (b) compare the dressed and bare phonon modes at two temperatures ($\sigma$ = 0.02 and 0.0057 Ry) obtained by solving the full $9\times 9$ dynamical matrix. Modes 1, 4, 5, and 9 have no electron-phonon coupling, so that $\omega_{\bf q}=\Omega_{\bf q}$ at eac temperature, although these modes still exhibit a weak $T$-dependence, panel (c). Other modes display stronger $T$-dependence, especially the longitudinal modes. We will see below that most of the $T$ dependence of the transverse modes is associated with anticrossing related weight transfer from longitudinal to transverse modes.
@@
Figure 5(d) provides evidence of a Kohn ladder related to the softening of modes 8, 7, and 3. The transverse modes have a weaker $T$-dependence, mainly in the $q$-regions where the self-energy shows transfer of spectral weight from longitudinal bands, and both bands show regions of phonon hardening at low temperatures. Notably, mode 7 shows strong softening at higher $T$s, which is not as strong when it interacts with mode 6, where mode 3 starts to soften more rapidly.
VII A Model of a Kohn ladder

### v2

@@ 1.2 Calculating electron-phonon coupling
We will refer to $\Pi^{\prime}$ as an auxiliary self-energy since it has dimension (energy)2. Note that the overall phase $e^{i\varphi_{\boldsymbol{k}}}$ as a result of the gauge transformation of the Bloch eigenfunction $|\Psi_{\boldsymbol{k}}\rangle$ in Eq.(5) in the context of the gauge non-invariance of $\mathfrak{g}_{\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ in the single band limit does not affect the off-diagonal elements of $\Pi^{\prime}$.
@@
To understand the origin of phonon softening, we examine ab initio and many-body results based on Eqs. (2) and (7). This requires a modification of the QE+EPW results, since the DFPT already incorporates the main phonon renormalizations (LABEL:SM3). To calculate $\Pi^{\prime}$ in Eq.(8), we must invert the dynamical matrix to get a handle on the nature of the bare phonons. This involves two issues. (i) There is ambiguity in inverting the dynamical matrix, which we resolve in Sections A Model of a Kohn ladder and LABEL:SM3. And (ii) we need the off-diagonal self-energy elements, which are essential for the adiabatic term, even though these are unimportant for the non-adiabatic corrections; extraction of off-diagonal self-energies elements is discussed in LABEL:SM2.
1.3 What determines $\boldsymbol{Q}_{CDW}$?
@@
A debate has arisen as to whether $\boldsymbol{Q}_{CDW}$ is dominated by the susceptibility in Eq.(6) or the electron-phonon matrix elements in Eq.(3). However, there is another potential mechanism. In Eq.(4) , the pre-factor contains $1/{\omega_{\boldsymbol{q}}}$ in $|\mathfrak{g}|^{2}$. This, following Hedin38, can be replaced by $1/{\Omega_{\boldsymbol{q}}}$, which acts as a Stoner denominator and causes the diagonal element $\Pi_{\lambda\lambda}$ to diverge when $\Omega_{\bf q}\to 0$ for $\boldsymbol{q}=\boldsymbol{Q}_{CDW}$ as $T\to T_{CDW}$. A similar argument has been invoked for the CDW physics in the cuprates.63
@@
Fig. 2 presents a decomposition of the self-energy of all phonon modes in panel (a) along the $\Gamma\rightarrow M$ symmetry line. We first focus on the soft-phonon mode (dark green lines). $\Pi$ (panel (h)) shows a sharp, nearly divergent peak at the soft-mode $\boldsymbol{q}$-vector, which arises entirely from the $1/\Omega_{q}$-factor (panel (g)). In contrast, $\Pi^{\prime}=\omega\Pi$ (panel (d)) has a weak peak at the same $\boldsymbol{q}$-vector. This sharpening of a weak peak is a typical effect of a Stoner denominator. If the electron-phonon matrix element were $\boldsymbol{q}$-independent, then $\Pi^{\prime}$ would peak at the same $\boldsymbol{q}$-value as the susceptibility $\chi$ (dashed red curve in panel (b)), which is not the case. Similarly, if $\chi$ were constant, the peak would follow $G$ (green curve in panel (b)), but that is also not the case. The product $G\chi$ (panel (c)) comes closer, indicating that both $\chi$ and $G$ possess significant $\boldsymbol{q}$ and $\boldsymbol{k}$ dependence, such that the full integral over $\boldsymbol{k}$ correctly reproduces the self-energy peak in panel (d). The peaks in both $\Pi$ and $\Pi^{\prime}$ (solid-black vertical lines) are in reasonable agreement with the experimental value $\boldsymbol{Q}_{CDW}\simeq 2~{\overrightarrow{\Gamma M}}/3$ (black vertical dashed line). Note however that the off-diagonal terms in $\Pi^{\prime}$ [panel (f)] due to the phonon mode-mode coupling can lead to a small additional shift of the $\boldsymbol{Q}_{CDW}$.
@@
When we extend the preceding discussion to encompass the self-energy in the full Brillouin zone, we see that (Fig. 1 (b)) the largest self-energy peak falls along the $M\rightarrow K$-line. This sharp peak is associated with a weak nesting of the Fermi surface, see LABEL:SM2 for details. Insight can be gained by recalling that in many superconductors, the superconducting gap can be captured in the optical spectra only in the dirty limit where momentum conservation breaks down64. A similar mechanism may be at play with the second peak along the $M\rightarrow K$-line in our case and its absence in experiments. Note that Figure 1(b) is for the clean limit. To capture the dirty limit, however, one must incorporate the breaking of the momentum conservation due to random impurities, which requires averaging $G_{\lambda,\lambda^{\prime}}$ over $\boldsymbol{k}$, leading to the form of Eq.(7). Fig. LABEL:fig:6hh (b) (LABEL:SM2) shows that the averaging procedure smears out this weak nesting feature while preserving the peak along $\Gamma-M$ with a modest loss of intensity. A similar conclusion was reached in Ref. 100.
1.4 Coupling of bare phonon bands
@@
To extract the bare phonons, the divergent $1/\Omega_{\lambda,\boldsymbol{q}}$-factor in Fig. 2(g) must be replaced by $1/\omega_{\lambda,\boldsymbol{q}}$ (SM 3). Once this is done, we can go beyond the above single mode picture by analyzing the $T$-dependence of the diagonal self-energies in Figs. 1 (b) and (c). We discuss the resulting anticrossing phenomena by considering two distinct cases below.
1.4.1 Longitudinal-transverse anticrossing
@@
Figure 1 (a) shows that reducing the temperature (right to left) results in crossing of the modes 2 and 3 and of modes 6 and 7 along the $\Gamma-M$ line, see Fig. 2 for notation of phonon modes. Note that the optical modes 6 and 7 display at low temperatures an unusual Mobius topology along the $\Gamma\to M\to K\to\Gamma$ line: if one follows the dispersion of mode 6, starting from the left-hand side $\Gamma$ point, one ends up on mode 7 on the right-hand side. Repeating this process into the next $\Gamma$-point by using symmetry to jump to the left-hand side $\Gamma$-point, and starting on mode 7, we see that only after winding twice about this path do we return to the original point. This behavior can be understood by noting that the longitudinal and transverse modes are orthogonal, so that we normally expect them to cross. However, in an anisotropic material, eigenmodes are strictly longitudinal/transverse only along the high symmetry lines but this is not the case along a general direction in the Brillouin zone, so that the bands then must anti-cross and exchange wavefunction character. Since modes 6 and 7 only cross once along $\Gamma-M$, they must wind twice to go from $6\rightarrow 7\rightarrow 6$ before returning to the starting mode.
@@
A similar phenomenon is at play for modes 2 and 3 in Fig.1, although it is hard to see the crossing between these modes because they are acoustic modes with vanishing frequencies at $\Gamma$. However, the most dispersive mode near $\Gamma$ is longitudinal (L, numbered 3) at the left-hand side $\Gamma$-point. The resulting mode is numbered 2 after the $\Gamma\to M\to K\to\Gamma$ loop is closed. Similar winding numbers have been observed in the cuprates under pressure71, see also Ref. 65. At higher temperatures (right column in Fig.1) both anti-crossings disappear, suggesting that they are related to phonon softening and electron-phonon interactions. Notably, crossing of bands 6 and 7 is symmetry protected only along the $\Gamma M$ line, where the crossing point forms the tip of a phononic Dirac dispersion. Since this band crossing is driven by phonon softening, it constitutes an electron-phonon-coupling induced topological phase transition.
@@
Insight into the crossing of modes 6 and 7 is also obtained by looking at the self-energies. As $T$ increases, the longitudinal mode 7 hardens and stops overlapping the transverse mode. At the crossing point, a part of the self-energy, especially along $K$ to $\Gamma$ jumps from the lower to the higher (now longitudinal) mode. [The strong spectral weight is always associated with the upper modes near $K$, but due to the change of connectivity, the labels of the modes change.] Effects of anticrossing can, however, still be seen in Figs. 1(b,c), where the spectral weight of the upper mode decreases substantially below the anticrossing temperature, i.e., as $T$ is lowered, the modes start to anticross, and more longitudinal character transfers to the lower modes, along with the strong electron-phonon coupling.
@@
1.4.2 Longitudinal-longitudinal anticrossing
Here, the anticrossing of the three longitudinal modes is harder to see in the dispersions as there are no symmetry-allowed crossings. Kohn-anomaly related anticrossings may thus be taken as a demanding test of a phonon code. Although we have shown that QE captures mode-crossing effects for modes 6 and 7, let us now consider modes 3 and 8, which are the lowest and highest longitudinal modes. As expected, the self-energy of mode 3 decreases with increasing $T$. However, the self-energy of mode 8 increases with increasing $T$, and as Figs. 1(b,c) indicate, this is due to transfer of spectral weight from mode 3 (note the anomalous peak along the $MK$ line marked by the arrow in panel (b)), which shifts mainly to the transverse mode 7 (arrow in panel (c)), and is starting to appear in mode 8 at the highest temperature studied. These results unambiguously demonstrate that the mode softening starts in the highest longitudinal mode, and as temperature decreases, the spectral weight transfers to successively lower longitudinal modes via mode coupling, with further spectral weight transfer between nominally longitudinal and transverse modes, since they are not strictly orthogonal at arbitrary $\boldsymbol{k}$-points. In the remainder of this paper, we will explore the dynamical matrix in order to extract a coherent picture in terms of ‘bare’ phonon frequencies.
@@ 1.4.2 Longitudinal-longitudinal anticrossing
Figure 3: (color online) Diagonal elements of phonon self-energy. Smearing of (from left to right) 0.0057, 0.009, and 0.02 Ry is used. From top to bottom: (a) clean-limit and (b) dirty limit; projected in color on phonon bands (colors from dark to light denote weak to strong phonon self-energies in meV).
1.5 Dynamical matrix and bare phonon frequencies
We can obtain the bare frequencies $\omega_{\bf q}(T)$ at various temperatures by inverting the dynamical matrix. We would expect the bare frequencies to be approximately independent of $T$, but that is not the case, see LABEL:SM3 for details. Note that phonon anharmonicity9 is not implicit in the DFPT calculations used in this work but this can be accessed using stochastic self-consistent harmonic approximation66 implemented in a more general DFPT formulation 32, 33.
We must first settle on the form of the dynamical matrix. In SI-1, we show that, despite some ambiguity, the appropriate choice is:
@@ 1.5 Dynamical matrix and bare phonon frequencies
which provides a conventional matrix equation for the bare phonon frequencies in terms of the dressed ones. Fig. 4 shows the bare frequencies (solid lines) obtained by neglecting the off-diagonal terms in Eq.(10). Although phonon softening can be seen, the mode anticrossings, which are controlled by the off-diagonal terms are absent.
@@
Figure 4: (color online) Single mode calculations for the three longitudinal modes (modes 3, 6, and 8). Solution of Eq.(SI.4) for $\sigma$ =0.0565 Ry, comparing the dressed ($\Omega_{q}$, dashed lines) and bare frequencies ($\omega_{q}$, solid lines).
@@
To understand why $\Omega_{\textbf{q}}$ becomes soft, we rewrite the single mode solution of Eq.(1) as $\Omega_{\bf{q}}^{2}=\omega_{\bf{q}}^{2}(1-S_{\bf{q}})$, where $S_{\bf{q}}=-2\Pi^{\prime}(\bf{q},\Omega_{\bf{q}})/{\omega_{\textbf{q}}^{\text{2}}}$ is an effective Stoner factor21. The instability (soft phonon) occurs at that $\bf q$ for which $S_{\bf q}\rightarrow 1$ first. At $\bf Q_{CDW}$, the bare frequency is $\omega_{\bf Q_{CDW}}=\sqrt{-2\Pi^{\prime}(\bf Q_{CDW})}$. If $\omega_{\bf q}$ depends weakly on $\bf q$, then the soft mode occurs when $|\Pi^{\prime}(\bf q)|$ is at its maximum.
1.5.1 Multimode calculations
@@
Figure 5: (color online) (a) Dressed (blue dashed lines) and bare (red solid lines) phonons for $\sigma$ = 0.02 Ry smearing. (b) Same as (a) but for $\sigma$ = 0.0057 Ry. (c) Temperature dependence of the bare phonons for $\sigma$ = 0.02 Ry (orange solid lines) and 0.0057 Ry (green dashed lines). (d) Temperature dependence of the dressed phonons, with colors representing various phonon modes: 2 (orange), 3 (red), 6 (magenta), 7 (green), and 8 (blue). $\sigma$ = 0.0057 Ry (dashed lines) and 0.02 eV (dotted lines), while the solid lines ($\sigma$ = 0.02 Ry bare phonons) are our best estimate of the high-$T$ limit.
@@
Figs. 5(a) and (b) compare the dressed and bare phonon modes at two temperatures ($\sigma$ = 0.02 and 0.0057 Ry) obtained by solving the full $9\times 9$ dynamical matrix. Modes 1, 4, 5, and 9 have no electron-phonon coupling, so that $\omega_{\bf q}=\Omega_{\bf q}$ at both temperatures, although these modes still exhibit a weak $T$-dependence, panel (c). Other modes display stronger $T$-dependence, especially the longitudinal modes. We will see below that most of the $T$ dependence of the transverse modes is associated with anticrossing related weight transfer from longitudinal to transverse modes.
@@
Figure 5(d) provides evidence of the Kohn ladder related to the softening of modes 8, 7, and 3. Notably, mode 7 shows strong softening at higher $T$s, which is not as strong when it interacts with mode 6, where mode 3 starts to soften more rapidly. Transverse modes have a weaker $T$-dependence, mainly in the $q$-regions where the self-energy shows transfer of spectral weight from longitudinal bands, and both bands show regions of phonon hardening at low temperatures.
1.6 A Model of a Kohn ladder
For gaining a handle on Kohn-ladder formation, we consider a 3-band (bands 2, 6, and 8) model calculation starting at high temperature. Following the toy model in LABEL:SM1, we assume that all the self-energy is initially concentrated on the highest mode. Specifically, for the model self-energy $\Pi^{\prime}_{m}$, we assume off-diagonal elements $|\Pi^{\prime}_{m,ij}|^{2}=20$ ($i\neq j$), $\Pi^{\prime}_{m,33}=\Pi^{\prime}_{m,66}$=0, and the temperature dependent self-energy is a scaled version of $\Pi^{\prime}_{33}$ in Fig. 2(b), $\Pi^{\prime}_{m,88}=\alpha\Pi^{\prime}_{33}$, where $\alpha$ increases as $T$ decreases.